OpenStudy (anonymous):

if carmen uses 1/2 lb of lettuce, 1 1/4 lb of tomatoes, and 1 1/3 lb of cucumber to make a salad. how many pounds of produce does carmen use in all

OpenStudy (anonymous):

(1/2) = (6/12) (1 1/4) = (5/4) = (15/12) (1 1/3) = (4/3) = (16/12) (6/12) + (15/12) + (16/12) = (37/12)